• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ole Hansen

    Log entry time 14:09:01 on March 28, 2012

    Entry number 367640

    keyword=Still investigating septum issues (if any)

    Jack and Jixie have been looking at the septum history. Jixie did change the right septum current around 13:35, which explains the voltage step. The oscillations of the coil voltage are correlated with the input colling water temperature and seem to be acceptable to Jack. The spike in the right septum field at the time of the left septum trip is simply crosstalk.

    So all seems fine,except that the left septum tripped for unknown reasons. Resetting.